A mixture is prepared by dissolving 1.70 g of ammonia gas and 5.35 g of ammonium chloride in distilled water and making up the volume to 100 mL.

Kb= 1.75 x10-5 at the particular temperature for NH3(aq)

The pH of the mixture would be:

a) 2.38

b) 9.24

c) 4.76

d) 11.68

e) 10.45
Answer #1

first calculate Molarity of NH3 and NH4Cl

Molarity = (W/MW) (1000 / V in mL)

[NH3] = (1.70 / 17)(1000 / 100) = 1.0 M

[NH4Cl] = (5.35 / 53.49) (1000/100) = 1.0 M

mixture of NH3 and NH4Cl act as basic buffer

pOH = pKb + log [NH4Cl] / [NH3]

pKb = - log Kb = - log [1.75 x10-5] = 4.75

pOH = 4.75 + log [1]/ [1]

pOH = 4.76

pH = 14 - 4.76

pH = 9.24

answer = option b = 9.24